# Spatio-Temporal Graphical Counterfactuals: An Overview

### v1

@@ IV-A Ignorability and Back-door Criterion
In the view of statistics, the ignorability assumption (see Assumption 4) is equivalent in calculation to the causal sufficiency assumption (see Assumption 8) with a back-door criterion. The accurate definition of back-door criterion can be found in these papers and books [53, 54, 16, 24]. Here we present another from with respect to the TABLE I, defined as
Definition 2 (Back-door Criterion [53, 54, 16, 24]).
@@
In TABLE I, the covariate $X$ is said to satisfy back-door criterion, if a Bayesian network is built on all the variables relative to $X,Y,T$, and there is an ordered pair $T\rightarrow Y$ that satisfy
@@
which is the same in case of $do(T=0)$ [53, 54, 16, 24]. Then, if the causal sufficiency is satisfied, that is, no hidden variable can interfere $X,Y,T$, the expectation of potential outcomes in TABLE I can be calculated by
@@
IV-B Counterfactual Falsifiability
Back to the example in Fig. 3, if we query to a POM, “Observed $X_{1}=0.5,X_{2}=1,X_{3}=1.5$, if had done $X_{2}=2$, what would $X_{3}=?$”, we must ensure there are at least two observational samples, and they have the same values $X_{1}=0.5$ but different values in $X_{2}$. That is, one is $X_{2}=1$, and another one is $X_{2}=2$, to satisfy the positivity assumption (see Assumption 3). And then, due to the consistency assumption (see Assumption 2), we can infer the counterfactuals through matching the two samples exactly. But this is different in SCM, we can infer the counterfactuals through the Pearl’s causal ladder, even if we only have a single sample, as shown in Section III-A. Thus, how to falsify the counterfactuals?
Figure 4: A diagram of back-door path.
Figure 5: A diagram of data extrapolation.
It is a difficult question, because the counterfactuals are unobserved according to the definition (see “?” in TABLE I), and we need at least two samples for comparison. In general, we cannot back to the past to falsify it with a different treatment, and we cannot reproduce the same experiments completely at different times. Thus, the inferred counterfactuals in SCM cannot be falsified without the assumptions of positivity and consistency, and this view is rejected ambiguously in Pearl’s book [26].
Another explanation is that, without the positivity, but with the consistency only in POM, the counterfactuals can be extrapolated by regressive approaches, and they are falsifiable. For example, as shown in Fig. 5, if there are $T$ and $Y$ only, and the observational data $(T,Y)=\{(0,0.5),(1,1),(2,2.5),(3,2)\}$ is collected, the linear regressive model can be fitted as $Y=0.5T+0.5$. Then, we query “if had done $T=4$, what would $Y=?$”, the counterfactual answer is $Y=2.5$, and the answer can be falsified by the regressive function because it does not beyond the observation. Thus, the consistency assumption actually claims that the counterfactuals can be falsified by the observations. This is also true for SCM, because SCM also uses the regressive approaches to infer counterfactuals.
IV-C Direct and Indirect Causation
@@
Another approaches are to test the significance of $H_{0}:b^{\prime}-b$ [84, 85, 86]. Or, to test the significance of $H_{0}:ab=0$ [87, 88]. If the correct regressive models are built, the counterfactuals can be calculated, just like SCM in Fig. 2. Moreover, they are all regression-based approaches, that restrict the usage in large multivariate datasets, because one needs to test whether a covariate is a confounder or a mediator, and the test process would be time-consuming seriously.
@@
But on the contrary, the test process can be faster with Bayesian network in an SCM. Here, we provide an algorithm to realize forward counterfactual inference as shown in Algorithm 1, and provide an example to show the inference process as shown in Fig. 7.
@@
The approaches to realize the codes in lines 1 and 3 in Algorithm 1 can be various, e.g., some variants of depth-first-search algorithm and breadth-first-search algorithm [89]. Thus, their time complexities are both $\mathcal{O}(|\mathbf{V}|+|\mathbf{E}|)$, if a Bayesian network $\mathcal{G}=(\mathbf{V},\mathbf{E})$ is given, where $\mathbf{V}$ is the node set and $\mathbf{E}$ is the edge set. Meanwhile, if the computational cost of factorization and adjustment are overlooked, the time complexity of the code in line 9 in Algorithm 1 is $\mathcal{O}(|\mathbf{V}|)$, because we only need to traverse all mediators in each pathway forward, and this process can be performed in parallel. Thus, the total time complexity of Algorithm 1 is $\mathcal{O}(|\mathbf{V}|^{2}+|\mathbf{V}||\mathbf{E}|)$. Thus, if the network structure of the Bayesian network is large, and it can be discovered accurately, the efficiency of SCM to infer counterfactuals would be better with polynomial time complexity.
@@
8             For every back-door path, select a group of direct predecessors, $\mathbf{W}\subset\mathbf{V}$, and then, according to back-door criterion (see Definition 2),
@@
×$\times\prod_{p}\prod_{q}P(X_{j_{p_{q}}}|do(X),X\in\mathbf{Pa}(X_{j_{p_{q}}})\subseteq\{X_{j_{1}:j_{p}}\}),$
@@
V Spatio-Temporal Graphical Counterfactuals
Back to the examples of counterfactual questions at the beginning of this article, that is, if a different investment strategy had been implemented, would we obtain higher returns? And, in computer networks, what changes would occur in network load if a node’s configuration had never been changed? And, would someone still purchase the product even if they had never been shown the advertisement? In these real-world scenarios, interactive behaviors with lagged causal effects are widely-existing. POM cannot answer these questions, because POM does not allow the interactive behaviors between experimental units (see SUTVA in Assumption 1), even with some temporal quasi-experimental approaches (e.g., Regression Discontinuity Design [90, 91] and Differences in Differences [92, 93]). And this is the same to SCM, because the foundational Bayesian network is a DAG that also does not allow mutual connections between network nodes. Thus, to answer these counterfactual queries for interactive units and lagged causal effects, a concept of spatio-temporal graphical counterfactuals is proposed here.
V-A Spatio-Temporal Bayesian Networks
Before the buildings of SCM, actually, various causal graphical models have been proposed to model the spatio-temporal causality. As shown in Fig. 8, Temporal Bayesian Network [64, 94] are proposed to model the momentary causal dependency in a first-order Markov process. Further, Dynamical Bayesian Networks [64] are proposed to represent a combination of multiple Temporal Bayesian Networks. Full Time Graph [25] is proposed to summarize the temporal Bayesian dependency at full timestamps. Note that, Full Time Graph allows the instantaneous causal effects between two units (or variables) in a high-order Markov process, but it only allows strictly stationary causality, that is, the causal dependencies do not change with time. Moreover, Full Time Graph is proved to uniquely decomposed into multiple high-order Temporal Bayesian Networks, named UCN, if the instantaneous causal effects is not allowed [95]. The comparisons of these causal graphs are shown in TABLE II.
Figure 8: Diagrams of various causal graphs.
TABLE II: Comparisons of various causal graphs.
@@ V-A Spatio-Temporal Bayesian Networks
Thus, based on the characteristics of these causal graphs, Spatio-Temporal Bayesian Networks (STBNs) are defined as a group of DAGs with different timestamps, as shown in Fig. 8. In STBNs, intuitively, the nodes are not equivalent to the variables, but represent the state of the variables at some time steps. With the assumptions of causal Markov (see Assumption 6) and faithfulness (see Assumption 7), the joint distribution of all nodes can be factorized as
@@
where $X_{1},\dots,X_{n}$ are the variables, and the timestamps are in range of $10\sim T-$ 1. Theoretically, infinite time steps are allowed in STBNs, but it is usually finite in practice.
To guarantee the global and unique DAG in STBNs, a temporal assumption is introduced as
Assumption 9 (Temporal Assumption).
In STBNs, the cause node precedes or parallels the effect node in time.
This assumption does not allow the directed cause-effect pairs from now to the past, e.g., $X_{i,t-\tau_{1}}\rightarrow X_{j,t-\tau_{2}},\tau_{1}<\tau_{2}$. And this guarantees the functionally identifiable network structure without Markov equivalence class [25, 95, 96]. Moreover, the causal sufficiency assumption (see Assumption 8) is also needed to guarantee the unconfoundedness.
@@ V-B Nonstationarity of STBNs
For the case of time-varying variance, the causal identifiability is up to the mutually independent exogenous additive noises [96]. That is, if there is no causal connection between two additive noise nodes, the causation is identifiable. Thus, an effective approach is to relate the time-varying noises with a common timestamp, and the timestamps at all time steps are observable confounders $C_{t-\tau},\tau=0,\dots,T-1$ [103, 97, 104, 98]. As shown in Fig. 9, there are three causal pathways, that is, $C_{t-1}\rightarrow U_{4,t-1}\rightarrow X_{4,t-1}$, $C_{t-1}\rightarrow C_{t}\rightarrow U_{3,t}\rightarrow X_{3,t}$, and $C_{t-1}\rightarrow C_{t}\rightarrow U_{4,t}\rightarrow X_{4,t}$. Thus, if $C_{t}$ and $C_{t-1}$ are hidden, the spurious connection between $X_{4,t-1}$ and $X_{3,t}$ would be detected, like the dotted line in Fig. 9. However, $C_{t}$ and $C_{t-1}$ are usually known as the timestamp $t$ and $t-1$, thus, $C_{t}$ and $C_{t-1}$ block the fork paths between $X_{4,t-1},X_{3,t}$ and $X_{4,t}$. Then, the noise variables $U_{4,t-1},U_{3,t}$ and $U_{4,t}$ are mutually independent. Thus, it is causally identifiable. Note that, however, the premise to do so is that any hidden confounders in the case of time-varying variance can be written as smooth or non-smooth functions of timestamps [98].
Nonstationary causality is challenging but critical, especially for the case of frequently changing structures. To identify the nonstationary causality, one approach is to use sliding windows to identify causation within different time periods, as shown in Fig. 9. For each sliding window, it is assumed that the causal structure is invariant, and then, a group of samples would be input to identification algorithms, e.g., [105, 106, 107, 108, 95], to identify a invariant structure during this period. However, if the network changes are frequent, the samples used for identification during a short-term period would be few, but in contrast, we usually need sufficient samples to accept the significance of causal connections. Thus, another reasonable approach is to identify the change points first, and then, identify the causality with the corresponding periods [109, 110]. However, this still does not work in the extreme case of frequently changing, because the samples between two change points may be few. Moreover, there are some other approaches to view the nonstationary causality as a probabilistic normalized flow, and then use variational Bayesian inference to estimate it globally [111, 112, 113].
V-C STBNs and Complex networks

### v2

@@ 4.2 Ignorability and Back-door Criterion
In the view of statistics, the ignorability assumption (see Assumption 2.2) is equivalent in calculation to the causal sufficiency assumption (see Assumption 3.1) with a back-door criterion. The accurate definition of back-door criterion can be found in these papers and books [17, 18, 4, 5]. Here we present another from with respect to the Table 1, defined as
Definition 4.1 (Back-door Criterion [17, 18, 4, 5]).
@@
In Table 1, the covariate $X$ is said to satisfy back-door criterion, if a Bayesian network is built on all the variables relative to $X,Y,T$, and there is an ordered pair $T\rightarrow Y$ that satisfy
@@
Figure 4: A diagram of data extrapolation.
Figure 5: A diagram of back-door path.
@@
which is the same in case of $do(T=0)$ [17, 18, 4, 5]. Then, if the causal sufficiency is satisfied, that is, no hidden variable can interfere $X,Y,T$, the expectation of potential outcomes in Table 1 can be calculated by
@@
4.3 Direct and Indirect Causation
@@
Figure 7: A diagram of front-door adjustment [5]. Here, $T$ is a treatment variable, and $Y$ is a target variable. $X_{1}$ and $X_{2}$ are both covariates.
@@
Another approaches are to test the significance of $H_{0}:b^{\prime}-b=0$ [77, 78, 79]. Or, to test the significance of $H_{0}:ab=0$ [80, 81]. If the correct regressive models are built, the counterfactuals can be calculated, just like SCM in Fig. 2. Moreover, they are all regression-based approaches, that restrict the usage in large multivariate datasets, because one needs to test whether a covariate is a confounder or a mediator, and the test process would be time-consuming seriously.
@@
But on the contrary, the test process can be faster with Bayesian network in an SCM. Here, we propose an algorithm to realize forward counterfactual inference as shown in Algorithm 1. To clearly demonstrate the principle of Algorithm 1, we first introduce the front-door criterion as follows:
Definition 4.2 (Front-door Criterion [17, 18, 4, 5]).
The covariate $X$ is said to satisfy front-door criterion, if a Bayesian network is built on all the variables relative to $X,Y,T$, and there is an ordered pair $T\rightarrow Y$ that satisfy
  1. 1.
    $X$ intercepts all directed paths from $T$ to $Y$;
  2. 2.
    there is no back-door path from $T$ to $X$;
  3. 3.
    all back-door paths from $X$ to $Y$ are blocked by $T$.
Then, a classic example is provided to show the front-door adjustment through the theorem 3.4.1 in book [5], as also shown in Fig. 7 here. If given a treatment variable $T$ and a target variable $Y$, with two covariates $X_{1}$ and $X_{2}$. Then one can adjust the distribution $P(Y|do(T))$ with the front-door criterion as follows:
$P(Y|do(T))$ $=\sum_{x_{1}}P(X_{1}|do(T))P(Y|do(X_{1}))$ (11)
$=\sum_{x_{1}}P(X_{1}|T)\sum_{t^{\prime}}P(Y|X_{1},T^{\prime})P(T^{\prime})$
Thus, we can conclude from this equation:
  1. 1.
    If there are one or multiple ordered causal pathways from $T$ to $Y$ and there is at least one mediator on the ways, then the causal pathway can be decomposed into two causal pathways through that mediator, e.g., $T\rightarrow X_{1}$ and $X_{1}\rightarrow Y$ in Eq. (11).
  2. 2.
    If there is no mediator on all causal pathways, then adjustment using back-door criterion. Of course, if there is no back-door path, no need to adjust.
Thus, the front-door principle is non-atomic and decomposable. And with this more foundational principle, we propose the forward counterfactual algorithm to autonomously apply this adjustment principle, as presented in Algorithm 1. The algorithm can be conducted on arbitrary variable set $\{X_{1},\dots,X_{n}\}$ that satisfied causal sufficiency. And we also provide an example to show the inference process as shown in Fig. 8.
@@
In Algorithm 1, the approaches to realize the codes in lines 1 and 3 can be various, e.g., some variants of depth-first-search algorithm and breadth-first-search algorithm [82]. Thus, their time complexities are both $\mathcal{O}(|\mathbf{V}|+|\mathbf{E}|)$, if a Bayesian network $\mathcal{G}=(\mathbf{V},\mathbf{E})$ is given, where $\mathbf{V}$ is the node set and $\mathbf{E}$ is the edge set. Meanwhile, if the computational cost of factorization and adjustment are overlooked, the time complexity of the code in line 9 in Algorithm 1 is $\mathcal{O}(|\mathbf{V}|)$, because we only need to traverse all mediators in each pathway forward, and this process can be performed in parallel. Thus, the total time complexity of Algorithm 1 is $\mathcal{O}(|\mathbf{V}|^{2}+|\mathbf{V}||\mathbf{E}|)$. Thus, if the network structure of the Bayesian network is large, and it can be discovered accurately, the efficiency of SCM to infer counterfactuals would be better with polynomial time complexity.
5 Spatio-Temporal Graphical Counterfactuals
Then for the examples of counterfactual questions at the beginning of this article, that is, if a different investment strategy had been implemented, would we obtain higher returns? And, in computer networks, what changes would occur in network load if a node’s configuration had never been changed? And, would someone still purchase the product even if they had never been shown the advertisement? In these real-world scenarios, interactive behaviors with lagged causal effects are widely-existing. POM cannot answer these questions, because POM does not allow the interactive behaviors between experimental units (see SUTVA in Assumption 2.2), even with some temporal quasi-experimental approaches (e.g., Regression Discontinuity Design [83, 84] and Differences in Differences [85, 86]). And this is the same to SCM, because the foundational Bayesian network is a DAG that also does not allow mutual connections between network nodes. Thus, to answer these counterfactual queries for interactive units and lagged causal effects, a concept of spatio-temporal graphical counterfactuals is proposed here.
@@ 5 Spatio-Temporal Graphical Counterfactuals
8      For every back-door path, select a group of direct predecessors, $\mathbf{W}\subset\mathbf{V}$, and then, according to back-door criterion (see Definition 4.1),
@@
$\times\prod_{p}\prod_{q}P(X_{j_{p_{q}}}|do(X),X\in\mathbf{Pa}(X_{j_{p_{q}}})\subseteq\{X_{j_{1}:j_{p}}\}),$
@@
Table 2: Comparisons of various causal graphs.
@@
5.1 Spatio-Temporal Bayesian Networks
Before the buildings of SCM, actually, various causal graphical models have been proposed to model the spatio-temporal causality. As shown in Fig. 9, Temporal Bayesian Network [59, 87] are proposed to model the momentary causal dependency in a first-order Markov process. Further, Dynamical Bayesian Networks [59] are proposed to represent a combination of multiple Temporal Bayesian Networks. Full Time Graph [6] is proposed to summarize the temporal Bayesian dependency at full timestamps. Note that, Full Time Graph allows the instantaneous causal effects between two units (or variables) in a high-order Markov process, but it only allows strictly stationary causality, that is, the causal dependencies do not change with time. Moreover, Full Time Graph is proved to uniquely decomposed into multiple high-order Temporal Bayesian Networks, named UCN, if the instantaneous causal effects is not allowed [21]. The comparisons of these causal graphs are shown in Table 2.
Figure 9: Diagrams of various causal graphs.
Thus, based on the characteristics of these causal graphs, STBNs are defined as a group of DAGs with different timestamps, as shown in Fig. 9. In STBNs, intuitively, the nodes are not equivalent to the variables, but represent the state of the variables at some time steps. With the assumptions of causal Markov (see Assumption 5) and faithfulness (see Assumption 6), the joint distribution of all nodes can be factorized as
@@ 5.1 Spatio-Temporal Bayesian Networks
where $X_{1},\dots,X_{n}$ are the variables, and the timestamps are in range of $0\sim T-1$. Theoretically, infinite time steps are allowed in STBNs, but it is usually finite in practice.
To guarantee the global and unique DAG in STBNs, a temporal assumption is introduced as {assumption}[Temporal Assumption] In STBNs, the cause node precedes or parallels the effect node in time. This assumption does not allow the directed cause-effect pairs from now to the past, e.g., $X_{i,t-\tau_{1}}\rightarrow X_{j,t-\tau_{2}},\tau_{1}<\tau_{2}$. And this guarantees the functionally identifiable network structure without Markov equivalence class [6, 21, 88]. Moreover, the causal sufficiency assumption (see Assumption 3.1) is also needed to guarantee the unconfoundedness.
@@ 5.2 Nonstationarity of STBNs
For the case of time-varying variance, the causal identifiability is up to the mutually independent exogenous additive noises [88]. That is, if there is no causal connection between two additive noise nodes, the causation is identifiable. Thus, an effective approach is to relate the time-varying noises with a common timestamp, and the timestamps at all time steps are observable confounders $C_{t-\tau},\tau=0,\dots,T-1$ [95, 89, 96, 90]. As shown in Fig. 10, there are three causal pathways, that is, $C_{t-1}\rightarrow U_{4,t-1}\rightarrow X_{4,t-1}$, $C_{t-1}\rightarrow C_{t}\rightarrow U_{3,t}\rightarrow X_{3,t}$, and $C_{t-1}\rightarrow C_{t}\rightarrow U_{4,t}\rightarrow X_{4,t}$. Thus, if $C_{t}$ and $C_{t-1}$ are hidden, the spurious connection between $X_{4,t-1}$ and $X_{3,t}$ would be detected, like the dotted line in Fig. 10. However, $C_{t}$ and $C_{t-1}$ are usually known as the timestamp $t$ and t1t-1, thus, $C_{t}$ and $C_{t-1}$ block the fork paths between $X_{4,t-1},X_{3,t}$ and $X_{4,t}$. Then, the noise variables $U_{4,t-1},U_{3,t}$ and $U_{4,t}$ are mutually independent. Thus, it is causally identifiable. Note that, however, the premise to do so is that any hidden confounders in the case of time-varying variance can be written as smooth or non-smooth functions of timestamps [90].
Nonstationary causality is challenging but critical, especially for the case of frequently changing structures. To identify the nonstationary causality, one approach is to use sliding windows to identify causation within different time periods, as shown in Fig. 10. For each sliding window, it is assumed that the causal structure is invariant, and then, a group of samples would be input to identification algorithms, e.g., [97, 98, 99, 100, 21, 22], to identify a invariant structure during this period. However, if the network changes are frequent, the samples used for identification during a short-term period would be few, but in contrast, we usually need sufficient samples to accept the significance of causal connections. Thus, another reasonable approach is to identify the change points first, and then, identify the causality with the corresponding periods [101, 102]. However, this still does not work in the extreme case of frequently changing, because the samples between two change points may be few. Moreover, there are some other approaches to view the nonstationary causality as a probabilistic normalized flow, and then use variational Bayesian inference to estimate it globally [103, 104, 105].
5.3 STBNs and Complex networks
@@ 6 Conclusion
[…]
This work is supported by the National Key R&D Program of China under Grant No. 2022ZD0120003, the National Natural Science Foundation of China under Grant Nos. 62233004, 62273090, 52525204, 52572354, the Jiangsu Provincial Scientific Research Center of Applied Mathematics under Grant No. BK20233002, the US National Science Foundation under Grant No. 2047488 and the Rensselaer-IBM AI Research Collaboration.